2 people and 1 horse. 1 person rode ahead and tied up the horse and kept walking. The 2nd person would would reach the point where the horse was tied up and begin riding until they reached the 1st person. They repeated this until they got home. The horse travels 5 mph and the people 3 mph. How long would it take to make a 20 mile trip?
Answer by scott8148(6628)

choose an arbitrary "step" size and calculate the average speed

the 1st man rides the horse 3 miles ___ this takes 3/5 hour

the 2nd man reaches the horse in 1 hour ___ the 1st man has 2/5 hour head start

r * t = d ___ 3[t + (2/5)] = 5t ___ 6/5 = 2t ___ 3/5 = t

so it takes 8/5 hr to travel 6 mi , where the two men and the horse are together to start another "step"
___ the average speed is [6 / (8/5)] or 15/4 mph ___ this is constant ,REGARDLESS of step size

you can see from the calculations , that both men walk and ride an equal distance

for the total trip ___ t = d / r ___ t = 20 / (15/4) = 16/3